how Cobb-Douglas can the function be used to calculate the sources of growth
1
Expert's answer
2021-08-03T09:01:41-0400

The Cobb- Douglas production function measures returns to scale by the sum of its exponents:

  • If "\alpha+\beta=1" , returns to scale are constant.
  • If "\alpha+\beta>1," returns to scale are increasing.
  • If "\alpha+\beta<1", returns to scale are decreasing.

The exponents of labor and capital in the Cobb-Douglas production function give a measure of output elasticities of labor and capital respectively.